OpenStudy (jenniferjuice):

Given the function f(x) = 2(x + 4), find x if f(x) = 20.

OpenStudy (rmrjr22):

solve accordingly

OpenStudy (rmrjr22):

20=2x+4.......20-4 = 2x.......16=2x....x =?

OpenStudy (jenniferjuice):

so um...

OpenStudy (rmrjr22):

16/2 = x

OpenStudy (jenniferjuice):

8

OpenStudy (rmrjr22):

i lied.... 2x +8 = 20.......20-8=2x......12=2x....x=?

OpenStudy (jenniferjuice):

where in the world did 8 come from?

OpenStudy (rmrjr22):

distributive property... 2(x+4) = 2x+8

OpenStudy (jenniferjuice):

but arent you not supposed to distribute with a function?

OpenStudy (rmrjr22):

yes, the function equals both 20 and 2x+8, so you set them equal to each other to find x

OpenStudy (rmrjr22):

you want to get x all by its self